(4.1)Use the reduced row echelon form to determine

2 4 6

0 0 2

2 -1 5


(4.2)Use the reduced row echelon form to determine

1 2 -4

2 -3 1

0 0 2


1
Expert's answer
2021-05-31T15:17:18-0400

4.1--


"\begin{pmatrix}
2 & 4 & 6 \\
0 & 0 & 2 \\
2 & -1 & 5
\end{pmatrix}"


(i) "R_1 \rightarrow \dfrac{R_1}{2}"


"\Rightarrow" "\begin{pmatrix}
1 & 2 & 3 \\
0 & 0 & 2 \\
2 & -1 & 5
\end{pmatrix}"


(ii) "R_3\rightarrow R_3-2R_1"

"\Rightarrow" "\begin{pmatrix}
1 & 2 & 3 \\
0 & 0 & 2 \\
0 & -5 & -1
\end{pmatrix}"


(iii) Interchange 2 and 3 row

"\begin{pmatrix}
1 & 2 & 3 \\
0 & -5 & -1 \\
0 & 0 & 2
\end{pmatrix}"


(iv) "R_2 \rightarrow \dfrac{R_2}{-5}"


"\begin{pmatrix}
1 & 2 & 3 \\
0 & 1 & 0.2 \\
0 & 0 & 2
\end{pmatrix}"


(v) "R_3 \rightarrow \dfrac{R_3}{2}"


"\begin{pmatrix}
1 & 2 & 3 \\
0 & 1 & 0.2 \\
0 & 0 & 1
\end{pmatrix}"


This is the Row Echelon Form of the matrix.

Since there are 3 non -zero rows hence the Rank is = 3


4.2--


"\begin{pmatrix}
1 & 2 & -4 \\
2 & -3 & 1 \\
0 & 0 & 2
\end{pmatrix}"


(i) "R_2\rightarrow R_2 -2R_1"

"\Rightarrow" "\begin{pmatrix}
1 & 2 & -4 \\
0 & -7 & 9 \\
0 & 0 & 2
\end{pmatrix}"


(ii) "R_2 \rightarrow \dfrac{R_2}{-7}" "\begin{pmatrix}
1 & 2 & -4 \\
0 & -1 & -9/7 \\
0 & 0 & 2
\end{pmatrix}"

"\Rightarrow" "\begin{pmatrix}
1 & 2 & -4 \\
0 & -1 & -9/7 \\
0 & 0 & 2
\end{pmatrix}"


(iii) "R_3 \rightarrow \dfrac{R_3}{2}"

"\Rightarrow" "\begin{pmatrix}
1 & 2 & -4 \\
0 & -1 & -9/7 \\
0 & 0 & 1
\end{pmatrix}"


This is the Row Echelon Form of the matrix.

There are 3 non -zero rows hence the rank of the matrix = 3